A container weighs 82.500 g when empty. When filled with 77.30 mL of a liquid, the combined mass is 153.620 g. What is the density of the liquid?
1
Expert's answer
2018-02-06T04:58:45-0500
dm = 153.620g - 82.500g = 71.12g
dV = 77.30 ml
\rho = dm/dV = 71.12 / 77.30 = 0.92 g/ml